OpenStudy (anonymous):

the equation of the tangent line to the graph of f(x) at (5, 52) is y=mx+b for m=25 and b=? f(x) = 3x^2-5x+2

OpenStudy (amistre64):

m = f'(x) sooo y = (6(5) -5)(x-5)+52

OpenStudy (anonymous):

thanks, how did you get that?

OpenStudy (amistre64):

well, they already gave the slope so the derivative was just a throw in

OpenStudy (amistre64):

but given a point P(x,y) and a slope of m; the equation can be determined by: y = m(x-Px)+Py

OpenStudy (anonymous):

ahh, I see what I was doing wrong.

OpenStudy (amistre64):

where Py = the b part...

OpenStudy (amistre64):

you can get there quicker by: y = mx -mPx + Py maybe :)

OpenStudy (anonymous):

yeah, that worked. thanks again

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

how would you get the intercept for this one then? f(x)=\[\sqrt{6-x}\] m=-.25

OpenStudy (anonymous):

the slope is given by the derivative at x=2

OpenStudy (anonymous):

crap, not slope. I meant intercept.

OpenStudy (amistre64):

the intercept is rather vague; do you mean the x intercept of the tangent line to the curve at x=2?